OpenStudy (anonymous):

Need HELP with this ϕ(x)=∫2π0(cos2(x+y)+1/2)ϕ(y)dy,

OpenStudy (amtran_bus):

I would need help with that too.

OpenStudy (anonymous):

is it like this\[\Large \phi(x)=\int\limits _0^{2 \pi}(\cos^2(x+y)+1/2)\phi(y)dy\]

OpenStudy (anonymous):

and wat is the question

OpenStudy (anonymous):

Yes it is. Myproblem is to put the kernelna way than i can use sin(ex) and cos2x

OpenStudy (anonymous):

Sin2x and cos2x brecasse the expressivos than i use most of the integrais are 0 and i dont get the values do a and b and lambda

OpenStudy (anonymous):

http://math.stackexchange.com/questions/404889/urgent-help-please-integral-eigenvectors-and-eigenvalues In the bottom is the method I use to solve this problems, find the eigenvectors and eigenvalues